How many liters of a 20% acid solution must be mixed with a 50 % acid solution in order to obtain 12 liters of a 30% solution?
Answer by stanbon(75887)

How many liters of a 20% acid solution must be mixed with a 50 % acid solution in order to obtain 12 liters of a 30% solution?
------------
Equation:
acid + acid = acid
0.50x + 0.20(12-x) = 0.30*12
------------
50x + 20*12 - 20x = 30*12
--------
30x = 10*12
-------
x = (1/3)12 = 4 liters (amt. of 50% solution needed)
-------
12-x = 8 liters (amt. of 20% solution needed)
-------------
Cheers,
Stan H.
-------------